7. (10 pts. Let (w) In (31 10). Find all of the necond order partial derivatives.
Answer #1

anstanta of are ang 2 and Y . ayon ay © Crven; f(x,y)? In (32-44) . In partial derivative, the variable, with respect to whicAirAL :():是(); from ] 加 3岁少 ; | 2014) 「an (3) du 57/ g : ) (5) [on or [5 ] (24) | 73时y) -13: +(+) = ( ) , ГД»» 07 с 4 да гу DPE to Alexande Now; ( \ се, 3 + чу я (2) = = ( ) : [wn ©] от 2 (1193) - 14 1чу) 2and order partial derivates of f(x,y) = ln (1304) are - 12 ayan Home to Perut 732 Tentang presen te